# OBLR-PO: A Theoretical Framework for Stable Reinforcement Learning

### v1

@@ Appendix A Proofs for Section 4
Recall that the noise term is defined as
@@
and the covariance matrix $\boldsymbol{H}(\theta)$ is given by
$\boldsymbol{H}(\theta)=\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\left(F(q,o)-b_{\theta}(q)\right)\right].$ (46)
@@
The covariance matrix of $\xi(\theta)$ is given by
$\text{Cov}[\xi(\theta)]=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)$ (47)
@@
$\text{Cov}[\xi(\theta)]$ $=\text{Cov}\left[\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right]-\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o)]\right]$ (48)
$=\frac{1}{N_{t}^{2}G_{t}^{2}}\sum_{j=1}^{N_{t}}\sum_{i=1}^{G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right]$ (49)
$=\frac{1}{N_{t}G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o)\right]$ (50)
$=\frac{1}{N_{t}G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\right]$ (51)
$=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta).$ (52)
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (53)
@@
$\|\nabla_{\theta^{\prime}}[\log f(x,\theta^{\prime})]-\nabla_{\theta}[\log f(x,\theta)]\|_{2}\leq L\|\theta^{\prime}-\theta\|_{2},$ (55)
@@
$\int\|\nabla_{\theta^{\prime}}[\pi_{\theta^{\prime}}(o|q)F(q,o)]-\nabla_{\theta}[\pi_{\theta}(o|q)F(q,o)]\|\mathop{}\!\mathrm{d}o\leq(BL+B^{2}M)\|\theta^{\prime}-\theta\|_{2}.$ (70)
@@
$\mathbb{E}[\mathscr{L}(\theta_{t+1})]\leq\mathbb{E}[\mathscr{L}(\theta_{t})]-\eta_{t}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{\eta_{t}^{2}(BL+B^{2}M)}{2}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (75)
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (76)
@@
$\eta_{t}=\frac{1}{BL+B^{2}M}\cdot\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}=\frac{1}{BL+B^{2}M}\cdot\frac{N_{t}G_{t}\operatorname{SNR}(\theta_{t})}{1+N_{t}G_{t}\operatorname{SNR}(\theta_{t})}.$ (77)
@@
$\operatorname{SNR}(\theta)=\frac{\|\nabla_{\theta}[J(\theta)]\|_{2}^{2}}{\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]}=\frac{\|\nabla_{\theta}\mathscr{L}(\theta)\|_{2}^{2}}{\operatorname{tr}(\boldsymbol{H}(\theta))}.$ (78)
@@
$\eta_{t}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\eta_{t}^{2}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{2(BL+B^{2}M)\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\right)}.$ (79)
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]$ $\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)$ (80)
$=\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}\cdot\frac{1}{2(BL+B^{2}M)}.$ (81)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right),$ (82)
@@
$\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}$ $\leq 2(BL+B^{2}M)\cdot\left(\mathbb{E}[\mathscr{L}(\theta_{0})]-\mathbb{E}[\mathscr{L}(\theta_{T})]\right)$ (88)
@@
$\left(\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}\right)$ $\cdot\left(\sum_{t=0}^{T-1}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)\right)$ (90)
$\geq\left(\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}$ (91)
@@
$\frac{\left(\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)}=O(1).$ (92)
@@
$1\gtrsim\frac{\left(\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)}\geq\frac{\left(\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)+4TB^{2}M}.$ (93)
@@
$\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\lesssim\frac{1+\sqrt{1+16TB^{2}M}}{2}\leq 1+\sqrt{4TB^{2}M}\lesssim\sqrt{T}.$ (94)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right)$ (95)
@@
As $\operatorname{tr}(\mathbf{H}(\theta))$ can be expressed as
$\operatorname{tr}(\mathbf{H}(\theta))$ $=\|\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\|_{2}^{2}$ (97)

### v2

@@ Appendix A Proofs for Section 4
Recall that the noise term is
@@
and we use $\boldsymbol{H}(\theta)$ and $\boldsymbol{C}(\theta)$ to denote the (single-sample) variance and the cross-sample covariance, respectively:
$\boldsymbol{H}(\theta)=\operatorname{Var}\!\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big)\right],$ (47)
$\boldsymbol{C}(\theta)=\operatorname{Cov}\!\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big),\,\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\big(F(q,o^{\prime})-b_{\theta}(q)\big)\right].$ (48)
@@
The variance matrix of $\xi(\theta)$ is given by
$\text{Var}[\xi(\theta)]=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)+\frac{G_{t}-1}{N_{t}G_{t}}\boldsymbol{C}(\theta).$ (49)
@@
$\text{Var}[\xi(\theta)]$ $=\text{Var}\left[\frac{1}{N_{t}}\sum_{j=1}^{N_{t}}\frac{1}{G_{t}}\sum_{i=1}^{G_{t}}\left[\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right]-\mathbb{E}[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o)]\right]$ (50)
$=\frac{1}{N_{t}^{2}G_{t}^{2}}\sum_{j=1}^{N_{t}}\text{Var}\left[\sum_{i=1}^{G_{t}}\nabla_{\theta}\log\pi_{\theta}(o_{i,j}|q_{j})\hat{A}^{\mathrm{PO}}(q_{j},o_{i,j})\right]$ (51)
$=\frac{1}{N_{t}G_{t}}\text{Var}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o)\right]$ (52)
$+\frac{G_{t}-1}{N_{t}G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)\hat{A}^{\mathrm{PO}}(q,o),\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\hat{A}^{\mathrm{PO}}(q,o^{\prime})\right]$ (53)
$=\frac{1}{N_{t}G_{t}}\text{Var}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\right]$ (54)
$+\frac{G_{t}-1}{N_{t}G_{t}}\text{Cov}\left[\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q)),\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)(F(q,o^{\prime})-b_{\theta}(q))\right]$ (55)
$=\frac{1}{N_{t}G_{t}}\boldsymbol{H}(\theta)+\frac{G_{t}-1}{N_{t}G_{t}}\boldsymbol{C}(\theta).$ (56)
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (57)
@@
$\|\nabla_{\theta}[\log f(x,\theta^{\prime})]-\nabla_{\theta}[\log f(x,\theta)]\|_{2}\leq L\|\theta^{\prime}-\theta\|_{2},$ (59)
@@
$\int\|\nabla_{\theta}[\pi_{\theta^{\prime}}(o|q)F(q,o)]-\nabla_{\theta}[\pi_{\theta}(o|q)F(q,o)]\|\mathop{}\!\mathrm{d}o\leq(BL+B^{2}M)\|\theta^{\prime}-\theta\|_{2}.$ (74)
@@
$\mathbb{E}[\mathscr{L}(\theta_{t+1})]\leq\mathbb{E}[\mathscr{L}(\theta_{t})]-\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{\eta_{t}^{2}(BL+B^{2}M)}{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))\right).$ (79)
It remains to control $\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))$. Recall that $\xi(\theta_{t})$ is an average of $N_{t}$ independent queries, each containing $G_{t}$ samples $o_{1},\dots,o_{G_{t}}$ drawn under the same $q$, so the variance decomposes into a within-sample term and a cross-sample term:
$\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))=\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))+\frac{G_{t}-1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{C}(\theta_{t})).$ (80)
Let $X:=\nabla_{\theta}\log\pi_{\theta}(o|q)\big(F(q,o)-b_{\theta}(q)\big)$ and $Y:=\nabla_{\theta}\log\pi_{\theta}(o^{\prime}|q)\big(F(q,o^{\prime})-b_{\theta}(q)\big)$, where $o$ and $o^{\prime}$ are distinct samples under the same $q$. For any unit vector $u$, Cauchy–Schwarz gives
$u^{\top}\operatorname{Cov}(X,Y)u=\operatorname{Cov}(u^{\top}X,\,u^{\top}Y)\leq\sqrt{\operatorname{Var}(u^{\top}X)\operatorname{Var}(u^{\top}Y)}=\operatorname{Var}(u^{\top}X),$ (81)
where the last equality uses that $X$ and $Y$ are identically distributed (given $q$). Hence $\operatorname{Cov}(X,Y)\preceq\operatorname{Var}(X)$, implying $\operatorname{tr}(\boldsymbol{C}(\theta_{t}))\leq\operatorname{tr}(\boldsymbol{H}(\theta_{t}))$. Therefore,
$\operatorname{tr}(\operatorname{Var}(\xi(\theta_{t})))\leq\frac{1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))+\frac{G_{t}-1}{N_{t}G_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))=\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t})).$ (82)
Thus, we have
$\mathbb{E}[\mathscr{L}(\theta_{t+1})]\leq\mathbb{E}[\mathscr{L}(\theta_{t})]-\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{\eta_{t}^{2}(BL+B^{2}M)}{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (83)
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$ (84)
@@
$\eta_{t}=\frac{1}{BL+B^{2}M}\cdot\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}=\frac{1}{BL+B^{2}M}\cdot\frac{N_{t}\operatorname{SNR}(\theta_{t})}{1+N_{t}\operatorname{SNR}(\theta_{t})}.$ (85)
@@
$\operatorname{SNR}(\theta)=\frac{\mathbb{E}\|\nabla_{\theta}[J(\theta)]\|_{2}^{2}}{\mathbb{E}[\|\nabla_{\theta}\log\pi_{\theta}(o|q)\,\hat{A}^{\mathrm{PO}}(q,o)-\nabla_{\theta}[J(\theta)]\|_{2}^{2}]}=\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta)\|_{2}^{2}}{\operatorname{tr}(\boldsymbol{H}(\theta))}.$ (86)
@@
$\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right).$
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{2(BL+B^{2}M)\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\text{tr}(\boldsymbol{H}(\theta_{t}))\right)}.$ (87)
@@
$\mathbb{E}[\mathscr{L}(\theta_{T})]$ $\leq\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\eta_{t}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{BL+B^{2}M}{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)$ (88)
$=\mathbb{E}[\mathscr{L}(\theta_{0})]-\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}\cdot\frac{1}{2(BL+B^{2}M)}.$ (89)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right),$ (90)
@@
$\sum_{t=0}^{T-1}\frac{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}$ $\leq 2(BL+B^{2}M)\cdot\left(\mathbb{E}[\mathscr{L}(\theta_{0})]-\mathbb{E}[\mathscr{L}(\theta_{T})]\right)$ (96)
@@
$\left(\sum_{t=0}^{T-1}\frac{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{4}}{\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))}\right)$ $\cdot\left(\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)\right)$ (98)
$\geq\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}$ (99)
@@
$\frac{\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)}=O(1).$ (100)
@@
$1\gtrsim\frac{\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}+\frac{1}{N_{t}}\operatorname{tr}(\boldsymbol{H}(\theta_{t}))\right)}\geq\frac{\left(\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)^{2}}{\sum_{t=0}^{T-1}\left(\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\right)+4TB^{2}M}.$ (101)
@@
$\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}\lesssim\frac{1+\sqrt{1+16TB^{2}M}}{2}\leq 1+\sqrt{4TB^{2}M}\lesssim\sqrt{T}.$ (102)
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla_{\theta}\mathscr{L}(\theta_{t})\|_{2}^{2}=O\left(\frac{1}{\sqrt{T}}\right)$ (103)
@@
As $\operatorname{tr}(\boldsymbol{H}(\theta))$ can be expressed as
$\operatorname{tr}(\boldsymbol{H}(\theta))$ $=\|\nabla_{\theta}\log\pi_{\theta}(o|q)(F(q,o)-b_{\theta}(q))\|_{2}^{2}$ (105)